OpenStudy (anonymous):

A ladder leaning against a house. The distance between the base of the ladder to the house is 3 feet. The length of the ladder is 10 feet. In feet how far above the ground does the ladder touch the house?

OpenStudy (anonymous):

i really need help plz help me thank ya in advance

OpenStudy (raden):



OpenStudy (raden):

solve for x by using the pythagorean theorem ; x^2 = 10^2 - 3^2

OpenStudy (anonymous):

still confused :(

OpenStudy (raden):

x^2 = 10^2 - 3^2 x^2 = 100 - 9 x^2 = 91 x = sqrt(91) = ... (calculator)

OpenStudy (anonymous):

so it's 9.5?

OpenStudy (raden):

yeah, hat's nearest

OpenStudy (anonymous):

thank ya :)

OpenStudy (raden):

wlcm :)